OpenStudy (dls):

Trigo help..where did I go wrong? \LARGE If ~\sin^{-1}x+\cot^{-1}1/2=\frac{\pi}{2}~Find~x.

OpenStudy (dls):

\[\LARGE If ~\sin^{-1}x+\cot^{-1}\frac{1}{2}=\frac{\pi}{2}~Find~x.\]

OpenStudy (dls):

I did it like this.. \[\LARGE ~\cancel \sin^{-1}x+\cot^{-1}\frac{1}{2}=\cos^{-1}x+\cancel \sin^{-1}x\]

OpenStudy (dls):

 From here..cosx=1/sqrt(5)